By method variation of parameters solve the ODE(D^2+4)y=2 cosec2x?


1
Expert's answer
2021-07-30T08:14:30-0400

Solution;

The general solution of the equation is given by;

y=yp+yh

The equation has an auxiliary equation of the form;

P(m)=m2+4=0

m2=-4

m=-2i and m=2i

Hence the homogenous solutions is;

yh=C1cos(2x)+C2sin(2x)

Here;

y1=cos(2x)

y2=sin(2x)

X=2cosec(2x)

"W=\begin{vmatrix}
 cos(2x) & sin(2x)\\
 -2sin(2x) & 2cos(2x)
\end{vmatrix}"

det(W)=2(cos2(2x)+sin2(2x))=2(1)=2"\ne0"

For the particular solution;

Let yp=uy1+vy2

"u=-\int\frac{y_2X}{W}dx"

"u=-\int\frac{sin(2x)2cosec(2x)}{2}" ="-\int1dx"

u=-x

"v=\int\frac{y_1X}{W}dx"

"v=\int\frac{cos(2x)2cosec(2x)}{2}dx" ="\int\frac{cos(2x)}{sin(2x)}dx" ="\frac12ln(sin(2x))"

Hence;

yp=-xcos(2x)+"\frac12"ln(sin(2x))sin(2x)

Hence the required solution is ;

y=C1cos(2x)+C2sin(2x)-xcos(2x)+"\frac12"sin(2x)ln(sin(2x))